### v1

# Primal-Dual iLQR

[…]
mm[#1]
Primal-Dual iLQR
João Sousa-Pinto (Apple)
joaospinto@gmail.com
   Dominique Orban (Polytechnique Montréal)
dominique.orban@polymtl.ca
(December 2023)
I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual SQP problem. As our algorithm is a specialization of NPSQP (Gill et al. 1992), it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS (Giftthaler et al. 2018) and FDDP (Mastalli et al. 2020).
II Introduction
II-A Unconstrained Discrete-Time Optimal Control Problems
Unconstrained discrete-time optimal control problems are optimization problems of the form
$&\min_{x_{0},u_{0}\ldots,x_{N}}&\sum\limits_{i=0}^{N-1}g_{i}(x_{i},u_{i})+g_{N}(x_{N})\\ &\mbox{s.t.}&x_{0}=s_{0}\\ &&x_{i+1}=f_{i}(x_{i},u_{i}),\forall i\in\{0,\ldots,N-1\}.\end{aligned}$
Such optimization problems are ubiquitous in the fields of motion planning and controls.
II-B Related Work
Naïf applications of generic optimization methods to unconstrained discrete-time optimal control problems would require $O(N^{3}(n+m)^{3})$ operations for solving the linear systems posed at each iteration, where $N,n,m$ are respectively the number of stages, states, and controls of the problem. The first algorithm to improve on this was DDP (Mayne 1966), achieving complexity $O(N(n+m)^{3})$. Its convergence properties were studied in (Shoemaker and Liao 1991). iLQR (Todorov and Li 2004) can be seen as a version of DDP with coarser second derivative information (in particular, disregarding the Hessians of the dynamics), trading off local quadratic convergence for some added ease of implementation. Stagewise Newton (Dunn and Bertsekas 1989) exploits the fact that the computation of a Newton step on the “eliminated problem” (obtained by eliminating the variables $x_{i}$ in decreasing order of $i$ by plugging in the dynamics into the costs) can be reduced to solving an LQR problem, thereby achieving the same computational complexity as DDP.
DDP, iLQR, and Stagewise Newton are all single shooting methods. This makes them hard to warm start, as the state trajectory cannot be chosen independently of the control trajectory. This can result in more iterations being required for these methods to converge, as well as higher likelihoods of convergence to undesired local optima.
(Giftthaler et al. 2018) introduces a multiple shooting equivalent of iLQR, called GNMS, as well as an algorithm that combines single and multiple shooting, called iLQR-GNMS, with similar computational complexity to the methods discussed above. There are no available convergence results for (iLQR-)GNMS; at a minimum, a line search procedure or a filter method would have to be added for these algorithms to achieve global convergence. (Mastalli et al. 2020) introduces an alternative multiple shooting equivalent of iLQR, called FDDP, which adds a curvilinear line search that smoothly transitions between using the nonlinear dynamics in the forward pass for larger step sizes and using the affine dynamics for smaller step sizes. There are also no available convergence results for FDDP. In both cases, as second order derivatives of the dynamics are not considered, there cannot be local quadratic convergence.
Much work has also been done in handling discrete-time optimal control problems with stagewise constraints. ALTRO (Howell, Jackson, and Manchester 2019), FATROP (Vanroye et al. 2023), and acados (Verschueren et al. 2022) are likely the most competitive open source software packages for this application at the time of writing. ALTRO handles constraints via the augmented Lagrangian method, while FATROP uses an interior point method and acados uses an SQP approach. ForcesNLP (Zanelli et al.) is likely the most competitive competing commercial software package at the time of writing, and also relies on an interior point method.
(Singh, Slotine, and Sindhwani 2022) introduces a family of methods combining SQP and DDP, one of which using a similar approach to the one discussed in this paper (what the authors call the “open loop (OL)” variant), which seems to perform poorly in their benchmarks. At several points in their paper and code (available in the Trajax repository on GitHub), there are missing affine terms, for example in the linearized dynamics, which would likely explain the benchmark results.
At a high level and modulo technicalities, FATROP can be seen as a specialization of IPOPT (Wächter and Biegler 2006) with a linear system solver that handles LQR problems with extra stagewise affine constraints in time $O(N(n+m)^{3})$. Inequality constraints are handled by eliminating some of the block-rows of the modified KKT system, as originally done in (Rao, Wright, and Rawlings 1998). The modified KKT system will only have the extra stagewise affine constraints when the problem being solved has other stagewise equality constraints besides the dynamics. A similar algorithm that does not handle other general equality constraints had previously been presented in (Pavlov, Shames, and Manzie 2020). As compared to the specialization of FATROP to problems without stagewise equality constraints, our method has the advantage of not requiring second order corrections for fast local convergence. ForcesNLP also follows an interior point approach, but relies on a Cholesky factorizations instead of LQR decompositions for solving its linear systems.
On the other hand, the augmented Lagrangian technique employed by ALTRO yields linear systems that can be solved via the LQR algorithm, even in the presence of stagewise equality constraints. However, ALTRO has a slightly unorthodox way of doing multiple shooting: it modifies the dynamics $x_{n+1}=f_{n}(x_{n},u_{n})$ into $x_{n+1}=f_{n}(x_{n},u_{n})+e_{n}$ and adds the constraints $e_{n}=0$. This makes it possible to independently warm start the state trajectory (while breaking the non-dynamic $e_{n}=0$ constraint) while still using a single shooting method. This approach has the disadvantage of substantially expanding the control dimension.
Finally, the SQP approach followed by acados poses inequality-constrained quadratic subproblems, which are then solved by HPIPM (Frison and Diehl 2020), an efficient interior point QP solver that exploits the discrete-time optimal control problem structure. While it does not currently support stagewise equality constraints other than the dynamics, it could be extended to do so using a similar approach to (Vanroye et al. 2023). Moreover, due to not employing a globalization strategy, global convergence is not guaranteed. When the functions defining the problem are expensive to evaluate, SQP methods become time-competitive, as the higher number of linear system solves will be compensated by the lower number of evaluations of the nonlinear problem.
II-C Contributions
This paper presents a self-contained description of the specialization of NPSQP to unconstrained discrete-time optimal control problems. As compared to GNMS and FDDP, our method has the advantages of being known to be globally convergent and to not impede superlinear local convergence. Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
This framework can be extended to also handle constrained problems, using augmented Lagrangian (Howell, Jackson, and Manchester 2019), interior point (Vanroye et al. 2023), or SQP (Verschueren et al. 2022) methods. ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses). Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
III General Optimization Background
III-A First-order Optimality Conditions and KKT Systems
Given a constrained optimization problem of the form
$\min_{x}$ $g(x)$ (1)
s.t. $c(x)=0,$ (2)
for any local optimum $x$ meeting the standard linear independence constraint qualification (LICQ) constraints (i.e., for which the rows of $J(c)(x)$ are linearly independent), there exists a vector $\lambda$ (called the Lagrange/KKT multiplier) for which
$\begin{cases}\nabla g(x)+\nabla c(x)\lambda&=0\\ c(x)&=0.\end{cases}$ (3)
In other words, $(x,\lambda)$ is a critical point (but not necessarily a minimizer) of the Lagrangian function $\mathcal{L}(x,\lambda)$ defined as $g(x)+\lambda^{T}c(x)$.
III-B Sequential Quadratic Programming
The sequential quadratic programming (SQP) method for equality-constrained non-linear optimization problems consists of applying Newton’s method for finding zeros of $\nabla\mathcal{L}$, typically in combination with a line search mechanism, in order to ensure global convergence.
Below, let $A=J(c)(x)$, $l=\nabla_{x}\mathcal{L}(x,\lambda)$, and $d=c(x)$. Moreover, let $Q$ be a positive definite approximation of $\nabla_{xx}^{2}\mathcal{L}(x,\lambda)$. Such an approximation can be constructed, for example, via regularization. Alternatively, the terms involving $\nabla_{xx}^{2}c(x)$ can be dropped entirely, and $Q$ would consist purely of a positive definite approximation of $\nabla_{xx}g(x)$. The latter approach, of course, would result in local quadratic convergence being lost. For simplicity of notation, we henceforth drop the dependencies of these terms in $x$.
Each Newton step for finding a zero of $h(z)$ is obtained by solving $h^{\prime}(z)\Delta z=-h(z)$. If $h=\nabla\mathcal{L}$, this corresponds to solving
$\begin{pmatrix}Q&A^{T}\\ A&0\end{pmatrix}\begin{pmatrix}\Delta x\\ \Delta\lambda\end{pmatrix}=-\begin{pmatrix}l\\ d\end{pmatrix}.$ (10)
It can be shown (by a simple application of the first-order optimality conditions mentioned above) that $\Delta x$ is the solution of
$\min_{p}$ $\frac{1}{2}p^{T}Qp+l^{T}p$ (11)
s.t. $Ap+d=0.$ (12)
Similarly, it can also be shown that $\Delta\lambda$ is the corresponding Lagrange multiplier. As long as $Q$ is positive definite, this cost function is bounded from below. Depending on the sparsity pattern of this system, weaker requirements on $Q$ might also guarantee this.
Convergence is established when $\Delta x$ and $\Delta\lambda$ are sufficiently close to 00. The solution is feasible as long as $d=0$, but convergence to locally infeasible solutions is possible.
III-C Merit Functions
Merit functions offer a measure of progress towards the final solution and are typically used in combination with a line search procedure. Note that it would be inadequate to simply apply line search on the original objective, as SQP may (and often does) provide a search direction that locally increases that objective (usually trading that off for a reduction in the constraint violations). Filter methods, used in (Wächter and Biegler 2006), provide an alternative mechanism for measuring progress and determining whether the candidate step should be accepted, but this paper will not be using that approach.
Given a penalty parameter $\rho>0$, we define the $\ell_{2}$ merit function
$m_{\rho}(x,\lambda)=\mathcal{L}(x,\lambda)+\frac{\rho}{2}\lVert c(x)\rVert^{2}.$ (13)
Below, $\Delta x$ and $\Delta\lambda$ correspond to SQP step defined above. Noting that $\nabla_{\lambda}\mathcal{L}(x,\lambda)=c(x)=d$, it holds that
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&0\end{pmatrix})$ (14)
$=\Delta x^{T}\nabla_{x}m_{\rho}(x,\lambda)$
$=\Delta x^{T}\nabla_{x}\mathcal{L}(x,\lambda)+\rho\Delta x^{T}J(c)(x)^{T}c(x)$
$=-\Delta x^{T}Q\Delta x-\Delta x^{T}A^{T}\Delta\lambda+\rho\Delta x^{T}A^{T}d$
$=-\Delta x^{T}Q\Delta x+\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}\Delta\lambda-\rho\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}d$
$=-\Delta x^{T}Q\Delta x+d^{T}\Delta\lambda-\rho\lVert d\rVert^{2}$
and that
$D(m_{\rho};\begin{pmatrix}0&\Delta\lambda^{T}\end{pmatrix})$ $=\nabla_{\lambda}\mathcal{L}(x,\lambda)^{T}\Delta\lambda=d^{T}\Delta\lambda.$ (15)
Therefore:
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})$ (16)
$=$ $-\Delta x^{T}Q\Delta x+2d^{T}\Delta\lambda-\rho\lVert d\rVert^{2}.$
Note that, as $Q$ is positive definite, $-\Delta x^{T}Q\Delta x<0$ unless $\Delta x=0$. When $d$ is not acceptably close to 00, we suggest taking
$\rho=\frac{2\lVert\Delta\lambda\rVert}{\lVert d\rVert},$ (17)
which ensures that $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0$ (due to the Cauchy-Schwarz inequality). Otherwise, we suggest taking $\rho=0.01$.
Note that (Gill et al. 1992) suggests instead never decreasing $\rho$, and whenever an increase is required, increasing it to twice the minimal value required for making
$D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0.$ (19)
III-D Line Search
Having computed $\Delta x$, $\Delta\lambda$, and $\rho$, we wish to compute a step size $\alpha$ that results in an acceptable decrease in our merit function $m_{\rho}$. Typically, this is achieved by performing a backtracking line search. $\alpha$ starts as $1$ and decreases by a constant multiplicative factor (often $0.5$) every iteration, until
$m_{\rho}(x+\alpha\Delta x,\lambda+\alpha\Delta\lambda)<$ (20)
$m_{\rho}(x,\lambda)+k\alpha D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})$
is satisfied (where $k\in(0,1)$ is called the Armijo factor; typically, $10^{-4}$ is used). This is called the Armijo condition. This process is guaranteed to terminate as long as $g(x)$ and $c(x)$ are differentiable at $x$ and $D(m_{\rho};\begin{pmatrix}\Delta x^{T}&\Delta\lambda^{T}\end{pmatrix})<0$.
When $\Delta x=0$ and $d=0$ but $\Delta\lambda\neq 0$, we take a full step without conducting a line search.
Note that NPSQP (Gill et al. 1992) uses a slightly different line search method. Specifically, it enforces both Wolfe conditions, not only the Armijo condition.
IV Algorithm Derivation
In this section, we specialize the methods described above to the case of unconstrained discrete-time optimal control problems as needed.
We start by defining
$d=\begin{pmatrix}d_{0}\\ \cdots\\ d_{N}\end{pmatrix}=c(x)=\begin{pmatrix}s_{0}-x_{0}\\ f(x_{0},u_{0})-x_{1}\\ \cdots\\ f(x_{N-1},u_{N-1})-x_{N}\end{pmatrix}.$ (28)
The corresponding Lagrange/KKT multipliers are $\lambda_{0},\ldots,\lambda_{N}$ respectively. Letting $x=(x_{0},u_{0},\ldots,x_{N-1},u_{N-1},x_{N})$, $A_{i}=J_{x}(f_{i})(x_{i},u_{i})$ and $B_{i}=J_{u}(f_{i})(x_{i},u_{i})$,
$A=J(c)(x)=$ (29)
$\begin{pmatrix}-I&0&0&0&0&\cdots&0&0&0\\ A_{0}&B_{0}&-I&0&0&\cdots&0&0&0\\ 0&0&A_{1}&B_{1}&-I&\cdots&0&0&0\\ 0&0&0&0&\ddots&\ddots&\ddots&0&0\\ 0&0&0&0&0&\cdots&A_{N-1}&B_{N-1}&-I\end{pmatrix}.$
Note that the LICQ conditions are always met for matrices of this form, due to the presence of the $-I$ blocks, independently of the $A_{i}$ and $B_{i}$.
The SQP constraint $Ap+d=0$ becomes
$\begin{cases}\Delta x_{0}=d_{0}\\ \Delta x_{i+1}=A_{i}\Delta x_{i}+B_{i}\Delta u_{i}+d_{i},\forall i\in\{0,\ldots,N-1\}.\end{cases}$ (30)
Note that if $x_{0}$ is warm started as $s_{0}$ then $\Delta x_{0}$ will always be 00.
Similarly, noting that
$\mathcal{L}(x,\lambda)=$ $\lambda_{0}^{T}(s_{0}-x_{0})+\sum\limits_{i=0}^{N-1}g_{i}(x_{i},u_{i})+$ (31)
$\lambda_{i+1}^{T}(f_{i}(x_{i},u_{i})-x_{i+1})+g_{N}(x_{N}),$
it holds that
$l=\nabla_{x}\mathcal{L}(x,\lambda)=\begin{pmatrix}l_{0}\\ \vdots\\ l_{N}\end{pmatrix}$ (35)
where
$\begin{cases}l_{i}=\nabla g_{i}(x_{i},u_{i})+\nabla f_{i}(x_{i},u_{i})\lambda_{i+1}-\lambda_{i},\\ \forall i\in\{0,\ldots,N-1\}\\ l_{N}=\nabla g_{N}(x_{N})-\lambda_{N}\end{cases}$ (36)
and that
$Q=\nabla_{xx}\mathcal{L}(x,\lambda)=\begin{pmatrix}Q_{0}&&\\ &\ddots\\ &&Q_{N}\end{pmatrix},$ (40)
where
$\begin{cases}Q_{i}=\nabla^{2}g_{i}(x_{i},u_{i})+\nabla^{2}f_{i}(x_{i},u_{i})\lambda_{i+1},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}=\nabla^{2}g_{N}(x_{N}).\end{cases}$ (41)
The SQP cost $\frac{1}{2}p^{T}Qp+l^{T}p$ is simply $\sum\limits_{i=0}^{N}\frac{1}{2}p_{i}^{T}Q_{i}p_{i}+l_{i}^{T}p_{i}$, where $p_{i}=(\Delta x_{i},\Delta u_{i})$ for $i\in\{0,\ldots,N-1\}$ and $p_{N}=\Delta x_{N}$.
If exact Hessians are used, the matrices $Q_{i}$ need not be positive semi-definite. However, we require that positive definite approximations be used instead. It would also be acceptable to only require positive semi-definiteness, as long as the $\nabla_{u_{i}u_{i}}^{2}$ components of the $Q_{i}$ be positive definite. However, this would require checking that $\Delta x^{T}Q\Delta x>0$ at the end of the LQR solve. If this condition is not met, positive definite approximations would have to be used. Requiring positive definiteness in the first place avoids this complication.
In this case, the primal part of the SQP problem is an LQR problem, which can be efficiently solved.
In order to ensure that the matrices $Q_{i}$ are positive semi-definite, it may be helpful to maintain regularization parameters $\mu_{i}$ and always use $Q_{i}+\mu_{i}I$ instead of $Q_{i}$. If, during the LQR process, we detect that $Q_{i}+\mu_{i}I$ is not positive definite, $\mu_{i}$ can be updated by multiplying it by an updated factor $r>1$. When increasing $\mu_{i}$ is not required, we can instead update $\mu_{i}$ by dividing it by $r$. Minimum and maximum regularization parameters may also be established, to allow, when possible, $\mu_{i}$ to eventually be set to 00 if $Q_{i}$ is consistently positive definite, as well as to prevent exploring unreasonably high regularization parameters.
Next, we show how to efficiently compute the $\Delta\lambda$, i.e. how to solve the dual part of the SQP problem. Having determined $\Delta x$, this can be done by specializing $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$. Note that we discard the rows where $J(c)(x)^{T}$ contains any of the $B_{i}$ (as they are not useful for computing the $\Delta\lambda_{i}$), and keep the rows where it contains any of the $A_{i}$.
Letting $l_{i}^{x},l_{i}^{u},Q_{i}^{xx},Q_{i}^{xu},Q_{i}^{ux},Q_{i}^{uu}$ be such that
$l_{i}=\begin{pmatrix}l_{i}^{x}\\ l_{i}^{u}\end{pmatrix},Q_{i}=\begin{pmatrix}Q_{i}^{xx}&Q_{i}^{xu}\\ Q_{i}^{ux}&Q_{i}^{uu}\end{pmatrix},\forall i\in\{0,\ldots,N-1\},$ (46)
it holds that
$\begin{cases}Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}-\Delta\lambda_{i}+A_{i}^{T}\Delta\lambda_{i+1}=\\ -l_{i}^{x},\forall i\in\{0,\ldots,N-1\}\\ Q_{N}\Delta x_{N}-\Delta\lambda_{N}=-l_{N}\end{cases}\Rightarrow$ (47)
$\begin{cases}\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x},\\ \forall i\in\{0,\ldots,N-1\}\\ \Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}.\end{cases}$
Note that these equations can be used to recursively compute the $\Delta\lambda_{i}$, in decreasing order of $i$. We call this method the dual LQR backward pass.
Moreover, note that these $\Delta\lambda_{i}$ also satisfy the block-rows of $\nabla_{xx}\mathcal{L}(x,\lambda)\Delta x+J(c)(x)^{T}\Delta\lambda=-\nabla_{x}\mathcal{L}(x,\lambda)$ that we discarded, as a solution to the full Newton-KKT system must exist due to the LICQ conditions being satisfied, and as these $\Delta\lambda_{i}$ are the only solution that satisfies the block-rows we did not discard.
Using the methods described above, we can efficiently compute both $\Delta x$ and $\Delta\lambda$. The remaining parts of the optimization algorithm we describe here don’t require further specialization in the case of unconstrained discrete-time optimal control problems.
V Algorithm Overview
In this section, we provide some pseudocode for the proposed algorithm. For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
def build_lqr(problem, x, u, v):
    n = problem.num_states()
    m = problem.num_controls()
    N = problem.num_stages()
    for i in range(N):
        cost_grad = problem.cost(i)
            .gradient().eval(x[:,i], u[:,i])
        dyn_grad = problem.dynamics(i)
            .gradient().eval(x[:,i], u[:,i])
        A_i = dyn_grad.transpose()[:,:n]
        B_i = dyn_grad.transpose()[:,n:]
        if i == 0:
            d_i = problem.initial_state()
                - x[:,0]
        else:
            y = problem.dynamics(i-1)
                .eval(x[:,i-1], u[:,i-1])
            d_i =  y - x[:,i]
        l_i = cost_grad +
            dyn_grad * costates[:,i+1]
            - costates[:,i]
        cost_hess = problem.cost(i)
            .hessian().eval(x[:,i], u[:,i])
        dyn_hess = problem.dynamics(i)
            .hessians().eval(x[:,i], u[:,i])
        Q_i = cost_hess
            + dyn_hess * costates[:,i+1]
    l_N = problem.cost(N).gradient()
        .eval(x[:,N]) - costates[:,N]
    Q_N = problem.cost(N).hessian()
        .eval(x[:,N])
    return LQR(A_*, B_*, d_*, Q_*, l_*)
def lqr_dual_solve(lqr, dx, du):
    N = lqr.num_stages()
    dv[:,N] = Qxx * dx[:,N] + lx
    for i in reversed(range(N)):
        Qxx = lqr.Q(i).xx()
        Qxu = lqr.Q(i).xu()
        At = lqr.A(i).transpose()
        lx = lqr.l(i).x()
        dv[:,i] = Qxx * dx[:,i] + Qxu * du[:,i]
            + At * dv[:,i+1] + lx
    return dv
def merit_rho(d, dv):
    if d == 0:
        return 0.01
    return 2 * dv.norm() / d.norm()
def eval_merit(problem, x, u, v, rho):
    g = problem.total_cost().eval(x, u)
    d = problem.total_constraints().eval(x, u)
    return g + v.transpose() * d
        + 0.5 * rho * d.squared_norm()
def slope(lqr, dx, du, d, dv, rho):
    N = lqr.num_stages()
    slope = -rho * d.squared_norm()
    for i in range(N):
        dxu = stack(dx[:,i], du[:,i])
        slope += lqr.q(i).transpose() * dxu
        slope += dv[:, i].transpose() * lqr.d(i)
    slope += lqr.q(N).x().transpose() * dx[:, N]
    slope += dv[:, N].transpose() * lqr.d(N)
    return slope
def line_search(problem, lqr, x, u, v,
                dx, du, dv, rho):
    alpha = 1
    m_o = eval_merit(problem, x, u, v, rho)
    slope = slope(lqr, dx, d, dv, rho)
    while alpha > kMinStepSize:
        x_n = x + alpha * dx
        u_n = u + alpha * du
        v_n = v + alpha * dv
        m_n = eval_merit(
            problem, x_n, u_n, v_n, rho)
        if m_n < m_o + kArmijo * alpha * slope:
            return alpha
    raise LineSearchFailure()
def primal_dual_ilqr(problem, states_ws,
        controls_ws, costates_ws):
    x = states_ws
    u = controls_ws
    v = costates_ws
    for it in range(kMaxIterations):
        lqr = build_lqr(problem, x, u)
        dx, du = lqr.Solve()
        dv = lqr_dual_solve(lqr, dx, du)
        if dx == 0 and du == 0:
            if dv == 0:
                return x, u, v
            else:
                v += dv
                continue
        rho = merit_rho(lqr.d(), dv)
        alpha = line_search(
            problem, lqr, x, u,
            v, dx, du, dv, rho)
        x += alpha * dx
        u += alpha * du
        v += alpha * dv
    return x, u, v
VI Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step.
Iteration Objective $\|c\|^{2}$ $m_{\rho}^{\prime}$ $\alpha$
1 69.1912 9.173966407775879 -196341.22 1.0
2 38.890675 10.204395294189453 -282.203 0.5
3 21.515545 9.039937973022461 -99.97078 0.25
4 12.537336 3.5910232067108154 -37.347027 0.5
5 13.005388 1.256274938583374 -7.008366 0.5
6 11.193476 0.05883491784334183 -8.02231 1.0
7 10.798406 0.013367442414164543 -0.79570323 1.0
8 10.654287 0.0059180837124586105 -0.2289545 1.0
9 10.586153 0.001393172424286604 -0.10068051 1.0
10 10.561506 0.0004326050984673202 -0.03127028 1.0
11 10.545758 0.00018692569574341178 -0.01689804 1.0
12 10.535523 8.762026118347421e-05 -0.01067685 1.0
13 10.528571 4.414142313180491e-05 -0.00720222 1.0
14 10.523667 2.3347358364844695e-05 -0.00506186 1.0
15 10.520094 1.2880265785497613e-05 -0.00367544 1.0
16 10.517421 7.388789981632726e-06 -0.0027444 1.0
17 10.515375 4.351216375653166e-06 -0.00209233 1.0
18 10.513782 2.6609566248225747e-06 -0.00163243 1.0
19 10.512507 1.6749986571085174e-06 -0.00129717 1.0
20 10.511484 1.07657911030401e-06 -0.00104529 1.0
21 10.510648 7.084668709467223e-07 -0.00085156 1.0
22 10.509963 4.725115729797835e-07 -0.00069964 1.0
23 10.509394 3.217612913886114e-07 -0.00057881 1.0
24 10.50892 2.2059847992750292e-07 -0.00048149 1.0
25 10.508526 1.53070544683942e-07 -0.00040247 1.0
26 10.508192 1.0723919530164494e-07 -0.00033782 1.0
27 10.507912 7.579520655554006e-08 -0.00028446 1.0
28 10.507676 5.401408387228912e-08 -0.0002403 1.0
29 10.507476 3.848523988381203e-08 -0.00020355 1.0
30 10.507305 2.7843480410183474e-08 -0.00017285 1.0
31 10.50716 2.0009650825159042e-08 -0.00014714 1.0
32 10.507039 1.4696987626905411e-08 -0.00012551 1.0
33 10.506933 1.0591413968086272e-08 -0.00010727 1.0
VII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm. This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in (Gill et al. 1992).
Our algorithm can also be used to significantly speed up ALTRO, one of the state of the art open source software packages for solving constrained discrete-time optimal control problems, by removing the control dimension expansion occurring therein. Our method also removes the need to handle second order corrections, which would simplify and potentially speed up FATROP.